[figure: 9.7 cm]

INTERACTION BETWEEN TLS AND CONDUCTION ELECTRONS IN ACTINIDE ARSENOSELENIDES

T. Cichorek^{a,b}, R.Wawryk^{b}, A. Wojakowski^{b}, Z. Henkie^{b}, J. Custers^{a}, P. Gegenwart^{a}, F.Steglich^{a}, E.D. Bauer^{c} and M.B. Maple^{c} aMax Planck Institute for the Chemical Physics of Solids, Dresden, Germany

bInstitute of Low Temperature and Structure Research, Polish Academy of Sciences,

Wroclaw, Poland cDepartment of Physics, University of California, San Diego, USA

In spite of strong ferromagnetism of UAsSe (T_{C}=113 K), its uncommon low-T properties i.e., a pronounced upturn in the electrical resistivity ρ(T), are certainly of non-magnetic origin. ^{77}Se NMR studies performed on powdered UAsSe clearly pointed to a crystallographic misplacement in the anionic sublattice. This gave additional evidence of interaction between the conduction electrons and the two-level-system (TLS) centers. A qualitatively similar ρ(T) behavior has also been observed in diamagnetic ThAsSe. We report ρ(T, p ≤1.8GPa) results obtained on UAsSe and ThAsSe single crystals in the temperature range 1.5 - 300 K. Although a high hydrostatic pressure was applied, only a small reduction in the ρ(T, p=0) upturn was observed. Moreover, the pressure did not affect the temperature at which the upturn starts to develop. Both facts fit well with the theoretical prediction of the non-magnetic TLS Kondo model. Discussion will be extended to the thermodynamic properties of actinide arsenoselenides.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Tomasz Cichorek

Address for correspondence : Noethnizter Str. 40 D-01187 Dresden Germany

Email address : cichorek@cpfs.mpg.de